# Non-abelian Rees construction and pure motives

### v1

@@ 4.2.1. Discussion of the sign conjecture
By Deligne [18], $\bar{\mathcal{C}}$ is tensor equivalent to the category of super-representations of a super group scheme $G$ of finite type. On the other hand, by Jannsen [22], $\bar{\mathcal{C}}$ is semisimple. The classification of semisimple categories of super-representations shows that the only non-reductive factors of $G$, if any, are orthosymplectic groups $OSp(1,2n)$, cf. [48].
@@
S1: $G$ is reductive171717i.e. is classical, not super. The term ‘$OSp(1,2n)$ factor’ is made very precise in [48]., i.e. has no $OSp(1,2n)$ factor.
@@
The tensor category ${\rm{Rep}}\,OSp(1,2n)$ is ‘well-understood’: mysteriously, it is essentially the same as ${\rm{Rep}}\,SO(2n+1)$ [40]. This may be used concretely to tackle S1.
@@
Under S1, for any indecomposable object $\mathcal{V}$} of $\mathcal{C}$}, $\pi({\mathcal{V}})$ is annihilated by some exterior or symmetric power. Under S2, the corresponding exterior or symmetric power of $\mathcal{V}$} is 0, hence $\mathcal{C}$} is a Kimura-O’Sullivan category in the sense of [6, 9.2]; conversely, any Kimura-O’Sullivan category $\mathcal{C}$} satisfies S1, S2. Besides, S2 may be replaced by
S2’: the radical of $\mathcal{C}$} [6, 1.4] is a tensor ideal (hence equal to the maximal tensor ideal $\mathcal{N}$}).
@@ 4.2.2. Künneth-algebraic classes
Let us define a Künneth-algebraic class to be a cohomology class in $H(Y)$ of the form $pr^{YZ}_{Y\ast}(\alpha\cup\pi_{YZ}^{i}\beta)$, where $\alpha$ and $\beta$ are algebraic cohomology classes on $Y\times Z$, and $\pi_{YZ}^{i}$ is some Künneth projector on $Y\times Z$. The same calculation as in [1, 2.1] shows that Künneth-algebraic classes and correspondences are stable under the usual operations.
They are in fact motivated classes, but closer to algebraic classes. In particular, they are almost algebraic (in Tate’s terminology), i.e. they specialize to algebraic classes by reduction over finite fields whenever $Y$ and $Z$ have good reduction (by [25]).
Using Künneth-algebraic correspondences, one can build tensor categories of ‘Künneth motives’ by imitation of Grothendieck’s procedure. They become even tensor categories after changing the commutativity constraint. The considerations of 4.1 then apply unconditionally to this context.
@@ 4.3. Specialization of motives
Taking for $\omega$ the fiber functor induced by etale $\ell$-adic cohomology, when $k$ is the fraction field of a discrete valuation ring with residue field of characteristic $\neq\ell$, there is a faithful tensor ‘reduction functor’ of motives with good reduction181818cf. [1, 9.4] in the context of motivated classes.. It is not known whether this functor preserves radicals (i.e. numerical equivalence), neither for Grothendieck motives nor for their ‘Künneth’ analogs. But Theorem 3.2.1 a) shows that there is an equivariant morphism between the corresponding quasi-homogeneous varieties which maps the open orbit to the open orbit.
@@ 4.4. An illustration: compact abelian pencils
There is a close connection between the Hodge conjecture for complex abelian varieties and the standard conjecture D for the total space $f:A\to S$ of a pencil of abelian varieties ‘of the same Hodge type’ parametrized by a smooth projective curve - which lies behind to the theorem that Hodge cycles on complex abelian varieties are motivated [1]. One may try to use the non-abelian Rees construction to improve on this result and show, in some cases, that some cohomology classes on the general fibers $A_{s}$ are algebraic (or at least Künneth algebraic).
@@
Note on Künneth projectors. It is not even known that $A$ satisfies the sign conjecture. However, in the process of showing that certain cohomology classes on $A_{s}$ are algebraic, one is only interested in certain direct factors $N$ of the relative motive $h(A/S)(i)$ (with realization $H(N)\subset Rf^{2i}_{\ast}K(i)$), and with the corresponding direct factor $M\subset h(A)(i)$ (with realization $H(M)\subset H(S,H(N))$), cf. [44, 6.1][31]. Moreover, one may replace $S$ by a finite etale cover $S^{\prime}$. In particular, if $H(N)$ is trivial as a local system, one may replace $S$ by a double cover $S^{\prime}$ and look at the action of the involution of $S^{\prime}$ on $H(S,H(N))$. On $H^{+}(S^{\prime},H(N^{\prime}))$, this involution acts trivially, but has no nonzero fixed point in $H^{1}(S^{\prime},H(N^{\prime}))$, hence represents the even Künneth projector on the pull-back $M^{\prime}\subset h(A_{S^{\prime}})$.
4.4.3.
However, separating $H^{0}(S,H(N))$ from $H^{2}(S,H(N))$ remains elusive. For simplicity, we shall rather work with Künneth motives, consider the direct factor $M^{0}$ of $h(A)(i)$ with realization $H^{0}(S,H(N))$, and ${\mathcal{C}}=\langle M^{0},N_{s}\rangle$. The pull-back by the inclusion of $s$ in $S$ gives a morphism of motives in $\mathcal{C}$}
@@
This is an isomorphism on realizations, and also in the category ${\mathcal{C}}_{tann}={\mathcal{C}}_{mot}$, but maybe not in $\mathcal{C}$}, and might even become the zero map in $\bar{\mathcal{C}}$ if Conjecture D fails.
Instead of taking a general point $s$, one may sometimes find a point $t\in S$ for which $H(N_{t})$ consists of algebraic classes, and consider
@@
In this case, $N_{s}$ is isomorphic to ${\bf 1}^{m}$ in ${\mathcal{C}}_{tann}$.
One may take $K=\mathbb{R}$ or $\mathbb{C}$ for instance. The non-abelian Rees construction attaches to $\mathcal{C}$} (resp. ${\mathcal{C}}_{1}=\langle M^{0}\rangle,{\mathcal{C}}_{2}=\langle N_{s}\rangle$) a quasi-homogeneous space $\rm X$ (resp. $\rm X_{1},X_{2}$) under the tannakian group $G$ of $\bar{\mathcal{C}}$ (resp. $G_{1}$ of $\bar{\mathcal{C}}_{1}$, $G_{2}$ of $\bar{\mathcal{C}}_{2}$). If ${\rm{char}}\,k=0$, Conjecture D is true for ${\mathcal{C}}_{2}$ and $\rm X_{2}$ is a point, so that all $G_{1}$-orbits in $\rm X$ are closed (Theorem 3.2.1 b)).
One has an equivariant map $\rm X\to\rm X_{1}$ and a section given by the specialization functor ${\mathcal{C}}\to{\mathcal{C}}_{1}$ induced by the specialization from $s$ to $t$ (Theorem 3.2.1 a)).
4.4.4. The example of Weil classes
@@
When $E^{+}=\mathbb{Q}$, such classes have recently been shown to be algebraic [30]. For other fields $E^{+}$, these motivated classes are not known to be algebraic (nor Künneth-algebraic).
@@
One can associate to Weil classes compact abelian pencils $A\to S$ as in [3]. For a general point $s\in S$, the various embeddings $E^{+}\hookrightarrow\mathbb{R}$ give rise to motives $N_{s}$ of rank $2$ with real coefficients (which split, after complexification, into the sum of a rank 1 motive $L_{s}$ and its complex-conjugate $\bar{L}_{s}\cong L^{\vee}_{s}$); to a motive $M^{0}$ of rank $2$ cut on $A$ (which also splits, after complexification, into the sum of two rank 1 motives $L,\bar{L}$); and to a morphism $\iota_{s}^{\ast}:M^{0}\to N_{s}$ (which splits after complexification into the sum of two morphisms $L\to L_{s},\bar{L}\to\bar{L}_{s}$).
If the Weil classes are (Künneth-)algebraic, the quasi-homogeneous space $\rm X$ is a point. If no tensor power of Weil classes happened to be (Künneth-)algebraic, one could easily conclude that $G_{1}$ must be the Deligne torus $\rm{Res}$$_{\mathbb{C}/\mathbb{R}}\,\mathbb{C}^{\ast}$, $G_{2}=U^{1}$, and $G=G_{1}\times G_{2}$ (via Goursat’s lemma). In this case, $\rm X_{1}$ would be a toric surface and $\rm X$ a toric variety of dimension $3$.
@@
Since $\rm X$ is defined over $\mathbb{R}$, complex conjugation, which acts on the character group $\mathbb{Z}^{3}$ of $G$ by exchanging the first two elements of a basis and multiplying the third one by 1-1, must stabilize the dual cone of $\rm X_{1}$ and $\rm X$ (cf. e.g. [32]). The dual cone of $\rm X_{1}$ thus contains the first quadrant and is symmetric with respect to the main diagonal; the dual cone of $\rm X$ contains the cone with four faces delimited by the vectors (1,0,1),(0,1,1),(1,0,1),(0,1,1)(1,0,1),(0,1,1),(1,0,-1),(0,1,-1).
@@
Going back to $\mathcal{C}$} through the non-abelian Rees dictionary, this argument of ‘reality’ shows (unconditionally) that, beside the morphisms $L\to L_{s},\bar{L}\to\bar{L}_{s}$ induced by $\iota_{s}^{\ast}$, there are also morphisms $L\to\bar{L}_{s},\bar{L}\to L_{s}$ in $\mathcal{C}$}. Can one give a geometric construction of them191919of course, Conjecture D would imply that these four motives are the unit motive $\bf 1$.?
@@ 4.5. Semisimplicity criteria and abelian motives
To finish, let us apply our semisimplicity criteria 3.4 to abelian motives. Künneth projectors are algebraic in this context [28].
4.5.1.
We start with an application of Criterion 3.4.1.
@@
This is well-known if char $k=0$ (Lieberman [28], without assumption on End $A$). Therefore, we may assume that $k$ is (a finitely generated) field of char. $p\neq\ell$. We take for $\omega$ the fiber functor induced by etale $\ell$-adic cohomology with coefficients in $\bar{\mathbb{Q}}_{\ell}$. The hypotheses of Criterion 3.4.1 are satisfied with $H^{\prime}=$ Zariski closure of the image of $Gal(k^{s}/k)$ in ${\rm{GL}}(H^{1}(A_{k^{s}},\bar{\mathbb{Q}}_{\ell}))$ (Tate-Zarhin). $\square$
4.5.2.
@@
Criterion 3.4.2 cannot be applied directly, since $\mathcal{C}$} contains the Tate twist and complex conjugation acts trivially on its characters of $\mathbb{G}_{m}$. We consider instead the category $\mathcal{C}$} of ungraded202020i.e. defined by algebraic correspondences of any degree. $\lambda$-adic homological motives on powers of $A$, with coefficients in $K$. This does not affect the issue of numerical equivalence versus $\ell$-adic homological equivalence.
@@
Note. This new proof of the Clozel-Deligne theorem, which uses very few features from the geometry of abelian varieties, extends to categories $\mathcal{C}$} of ‘CM motives’, not related a priori to abelian varieties: those for which the special numerical motivic group’ $G$ (the kernel of the map from the numerical motivic group to $\mathbb{G}_{m}$ associated to the subcategory of Tate motives in $\mathcal{C}$}) has the properties of Criterion 3.4.2.

### v2

@@ 4.2.1. Discussion of the sign conjecture
According to Deligne [18], $\bar{\mathcal{C}}$ is tensor equivalent to the category of super-representations of a super group scheme $G$ of finite type. On the other hand, According to Jannsen [23], $\bar{\mathcal{C}}$ is semisimple. The classification of semisimple categories of super-representations shows that the only non-reductive factors of $G$, if any, are orthosymplectic groups $OSp(1,2n)$, cf. [50].
@@
S1: $G$ is reductive141414i.e. is classical, not super. The term ‘$OSp(1,2n)$ factor’ is made very precise in [50]., i.e. has no $OSp(1,2n)$ factor.
@@
The tensor category ${\rm{Rep}}\,OSp(1,2n)$ is ‘well-understood’: it is essentially the same as ${\rm{Rep}}\,SO(2n+1)$, cf. [42]. This may be used concretely to tackle S1. For instance, if $\mathcal{C}$} is generated as an even tensor category by an indecomposable motive $M$ with coefficients in $K=\bar{K}$, such that its realization has dimension a power of $2$, then no factor $OSp(1,2n)$ occurs in $G$ since all irreducible representations of $OSp(1,2n)$ are odd-dimensional.
@@
Under S1, for any indecomposable object $\mathcal{V}$} of $\mathcal{C}$}, $\pi({\mathcal{V}})\in\bar{\mathcal{C}}$ is annihilated by some exterior or symmetric power. Under S2, the corresponding exterior or symmetric power of $\mathcal{V}$} is 0, hence $\mathcal{C}$} is a Kimura-O’Sullivan category in the sense of [6, 9.2]; conversely, any Kimura-O’Sullivan category $\mathcal{C}$} satisfies S1, S2. Besides, under S1, S2 is equivalent to
S2’: the radical of $\mathcal{C}$} [6, 1.4] is a tensor ideal (hence equal to the maximal tensor ideal $\mathcal{N}$}, cf. [6, 1.4.4.b), 7.1.6]).
@@ 4.2.2. Künneth-algebraic classes
Let us define a Künneth-algebraic class to be a cohomology class in $H(Y)$ of the form $pr^{YZ}_{Y\ast}(\alpha\cup\pi_{YZ}^{i}\beta)$, where $\alpha$ and $\beta$ are algebraic cohomology classes on $Y\times Z$, and $\pi_{YZ}^{i}$ is some Künneth projector on $Y\times Z$. The same calculation as in [1, 2.1] (replacing Lefschetz involutions by Künneth projectors) shows that Künneth-algebraic classes are stable under the usual operations: tensor product, inverse and direct image by projections $pr^{YZ\ast}_{Y},pr^{YZ}_{Y\ast}$, hence that Künneth-algebraic correspondences are stable under composition.
Since Künneth projectors are motivated, Künneth-algebraic classes are motivated classes, but they are closer to algebraic classes. In particular, they are almost algebraic (in Tate’s terminology), i.e. they specialize to algebraic classes by reduction over finite fields whenever $Y$ and $Z$ have good reduction (by [26]).
Using Künneth-algebraic correspondences, one can build tensor categories of ‘Künneth motives’ as in [1, 4.2] (replacing Lefschetz involutions by Künneth projectors). They become even tensor categories after changing the commutativity constraint. The considerations of 4.1 then apply unconditionally to this context.
Coming back to our notation $G,H$ for the tannakian groups of $\mathcal{C}$} and ${\mathcal{C}}_{tann}$ respectively, a cohomology class is $G$-invariant if and only if it is Künneth-algebraic, whereas (when ${\rm{char}}\,k=0$) it is $H$-invariant if and only if it is motivated.
@@ 4.3. Specialization of motives
Taking for $\omega$ the fiber functor induced by etale $\ell$-adic cohomology, when $k$ is the fraction field of a discrete valuation ring with residue field of characteristic $\neq\ell$, there is a faithful tensor ‘reduction functor’ ${\mathcal{C}}\to{\mathcal{C}}^{\prime}$ of homological motives with good reduction (cf. [1, 9.4] in the context of motivated classes). It is not known whether this functor preserves radicals (i.e. descend to a ‘reduction functor’ of numerical motives), neither for Grothendieck motives nor for their ‘Künneth analogs’.
But Theorem 3.2.1 a) shows that there is an equivariant morphism ${\rm X}^{\prime}\to{\rm X}$ between the corresponding affine quasi-homogeneous varieties, which maps the open orbit to the open orbit.
@@ 4.4. An illustration: compact abelian pencils
There is a close connection between the Hodge conjecture for complex abelian varieties and the standard conjecture D for the total space $f:A\to S$ of a pencil of abelian varieties ‘of the same Hodge type’ parametrized by a smooth projective curve - which lies behind to the theorem that Hodge cycles on complex abelian varieties are motivated [1]. One may try to use the non-abelian Rees construction to improve on this result and show, in some cases, that some Hodge classes on the general fibers $A_{s}$ are algebraic (or at least Künneth-algebraic).
@@
It is not even known if the total space $A$ of the abelian pencil satisfies the sign conjecture. In order to avoid such complications, which lie on the sidelines of our main viewpoint (thinking about pure homological motives as equivariant vector bundles), we shall work instead with Künneth motives, and consider the direct factor $M^{0}$ of $h(A)(i)$ with realization $H^{0}(S,H(N))$, and ${\mathcal{C}}=\langle M^{0},N_{s}\rangle$. The pull-back by the inclusion of $s$ in $S$ gives a morphism of motives in $\mathcal{C}$}
@@
This is an isomorphism on realizations, and already an isomorphism in the category ${\mathcal{C}}_{tann}={\mathcal{C}}_{mot}$ of ‘motivated motives’; but maybe not in the category $\mathcal{C}$} of ‘Künneth motives’, and it might even become the zero map in the quotient $\bar{\mathcal{C}}$ if Conjecture D fails.
Instead of taking a general point $s\in S$, one may sometimes find a point $t\in S$ for which $H(N_{t})$ consists of algebraic classes, and consider
@@
In this situation, $N_{s}$ is isomorphic to ${\bf 1}^{m}$ in ${\mathcal{C}}_{tann}$, since parallel transport preserves motivated classes [1, Th. 0.5], and the subgroup $H\subset G$ is trivial.
The non-abelian Rees construction attaches to $\mathcal{C}$} (resp. ${\mathcal{C}}_{1}=\langle M^{0}\rangle,\,{\mathcal{C}}_{2}=\langle N_{s}\rangle$) a quasi-homogeneous space $\rm X$ (resp. $\rm X_{1},X_{2}$) under the tannakian group $G$ of $\bar{\mathcal{C}}$ (resp. $G_{1}$ of $\bar{\mathcal{C}}_{1}$, $G_{2}$ of $\bar{\mathcal{C}}_{2}$).
If ${\rm{char}}\,k=0$, Conjecture D is true for ${\mathcal{C}}_{2}$, and $\rm X_{2}$ is a point, so that all $G_{1}$-orbits in $\rm X$ are closed (Theorem 3.2.1 b)).
On the other hand, one has a $G$-equivariant map $\rm X\to\rm X_{1}$, together with a $G$-equivariant section given by the specialization functor ${\mathcal{C}}\to{\mathcal{C}}_{1}$ induced by the specialization from $s$ to $t$ (Theorem 3.2.1 a) and §4.3).
4.5. The example of Weil classes
4.5.1.
@@ 4.5. The example of Weil classes
When $E^{+}=\mathbb{Q}$, such classes have recently been shown to be algebraic [32]. For other fields $E^{+}$, these motivated classes are not known to be algebraic, nor even Künneth-algebraic.
4.5.2.
@@
Abelian varieties carrying Weil classes of ‘the same type’ are parametrized by certain Shimura varieties of PEL type. As in [3, §1.5], one can construct compact abelian pencils $A\to S$ ‘of Weil type’ such that for a general point $s\in S$, the various embeddings $E^{+}\hookrightarrow\mathbb{R}$ give rise to a motive $N_{s}$ of rank $2$ with real coefficients (which splits, after complexification, into the sum of a rank 1 motive $L_{s}$ and its complex-conjugate $\bar{L}_{s}\cong L^{\vee}_{s}$); to a motive $M^{0}$ of rank $2$ cut on $A$ (which also splits, after complexification, into the sum of two rank 1 motives $L,\bar{L}$); and to a morphism $\iota_{s}^{\ast}:M^{0}\to N_{s}$, which splits after complexification into the sum of two morphisms $L\to L_{s},\bar{L}\to\bar{L}_{s}$, which are isomorphisms in ${\mathcal{C}}_{tann}={\mathcal{C}}_{mot}$.
Moreover, one may assume that there is a point $t\in S$ for which $H(N_{t})$ consists of algebraic classes, so that $H$ is trivial, cf. §4.4.2.
This easily implies that $G_{1}$ is contained in the Deligne torus $\mathbb{S}:={\rm{Res}}_{\mathbb{C}/\mathbb{R}}\,\mathbb{C}^{\ast}$, while $G_{2}$ is contained in $U^{1}$, and $G$ in the product $G_{1}\times G_{2}$ (and maps surjectively onto the factors). Remind that the character group of $\mathbb{S}$ is isomorphic to $\mathbb{Z}^{2}$ with complex conjugation permuting the two generators $z,\bar{z}$; closed subgroups of $\mathbb{S}$ correspond to quotients of this lattice invariant by exchanging the images of $z$ and $\bar{z}$.
4.5.3.
If the Weil classes (at a general point $s$) are algebraic (or Künneth-algebraic), then $G$ is trivial and the quasi-homogeneous space $\rm X$ is a point.
If the are not Künneth-algebraic, the only possibilities for $G_{1}$ are $\mathbb{S}$ or $\mathbb{G}_{m}$. Indeed, the other closed subgroups of $\mathbb{S}$ being extensions of a cyclic group $\mathbb{Z}/n\mathbb{Z}$ by $U^{1}$, one would have $(\wedge^{2}M^{0})^{n}\cong{\bf 1}\cong(\wedge^{2}N_{t})^{n}$, $\iota_{t}^{\ast}$ would be an isomorphism, which would imply in turn that $L,\bar{L},$ as well as $L_{s},\bar{L}_{s}$ would be $\cong\bf 1$, contrary to our assumption.
On the other hand, $G_{2}$ is either $U^{1}$ (if no tensor power of the Weil classes are Künneth-algebraic) or $\mathbb{Z}/n\mathbb{Z}$ for some $n>1$; and $G=G_{1}\times G_{2}$ by Goursat’s lemma.
@@
Let us look more closely at the case $G=\mathbb{S}\times U^{1}$ (the other cases are similar and simpler). In this case, $\rm X_{1}$ is a toric surface and $\rm X$ a toric variety of dimension $3$. Complex conjugation, which acts on the character group $\mathbb{Z}^{3}$ of $G$ by exchanging the first two elements of a basis and multiplying the third one by 1-1, must stabilize the dual cone of $\rm X_{1}$ and $\rm X$ (cf. e.g. [34]). The dual cone of $\rm X_{1}$ thus contains the first quadrant and is symmetric with respect to the main diagonal; the dual cone of $\rm X$ contains the cone with four faces delimited by the vectors (1,0,1),(0,1,1),(1,0,1),(0,1,1)(1,0,1),(0,1,1),(1,0,-1),(0,1,-1).
@@
Going back to $\mathcal{C}$} via the non-abelian Rees dictionary, this argument of ‘reality’ shows that, beside the morphisms $L\to L_{s},\bar{L}\to\bar{L}_{s}$ induced by $\iota_{s}^{\ast}$, there are also morphisms $L\to\bar{L}_{s},\bar{L}\to L_{s}$ in $\mathcal{C}$}. This shows in particular that some explicit correspondences (different from the Weil classes themselves) are Künneth-algebraic, hence come from algebraic classes on products. Can one give a geometric construction of them (independently of Conjecture D, which would imply that these four motives are the unit motive)?
4.5.4.
A similar analysis applies to other contexts, such as Kuga-Satake correspondences between K3 surfaces and certain abelian varieties defined over $k=\mathbb{C}$. They are motivated classes [1, 7.1]. In the proof, one considers a pencil $Y\to S$ of $K_{3}$ surfaces, the associated Kuga-Satake abelian pencil $A\to S$ (with $S$ a projective smooth curve), and the pencil $Y\times_{S}Y\times_{S}A$, and one cuts out appropriate motives.
One can investigate the catalog of possibilities for the pairs $(G,{\rm X})$ that appear in this context (say over $K=\bar{K}$), and find that there are all spherical. Denoting by $r$ the rank of the transcendental part of the $H^{2}$ of a general fiber $Y_{s}$, one expects (under Conjecture D) that $G=H$, a group of type $Spin(r)$.
A typical case for $(G,H)$ in the catalog of possibilities (independently of Conjecture D) is $G=\mathbb{G}_{m}\times SO(r)\times Spin(r)$, and $H=$ the diagonal subgroup $Spin(r)\subset SO(r)\times Spin(r)$. Note that it is the $\mathbb{G}_{m}$ factor that prevents in this case to apply Luna’s criterion as in §3.4.
@@ 4.6. Semisimplicity criteria and abelian motives
4.6.1.
Of course, it is interesting to see what the Galois-type correspondence discussed in this paper (Theorem 2.3.1), based on the non-abelian Rees construction, might say about Conjecture D in special cases.
To finish, we apply our two general semisimplicity criteria (§3.4) to abelian motives. Note that Künneth projectors are algebraic in this context [30].
4.6.2.
We start with an application of Criterion 3.4.1 (which relies on Luna’s criterion for closed orbits).
@@
This is well-known if char $k=0$ (Lieberman [30], without assumption on End $A$). Therefore, we may assume that $k$ is field of char. $p\neq\ell$, and even that $k$ is finitely generated over $\mathbb{F}_{p}$. We take for $\omega$ the fiber functor induced by etale $\ell$-adic cohomology with coefficients in $\bar{\mathbb{Q}}_{\ell}$. The hypotheses of Criterion 3.4.1 are satisfied with $H^{\prime}=$ Zariski closure of the image of $Gal(k^{s}/k)$ in ${\rm{GL}}(H^{1}(A_{k^{s}},\bar{\mathbb{Q}}_{\ell}))$ (Tate-Zarhin). $\square$
4.6.3.
@@
We end with an application of Criterion 3.4.2 (which relies on real descent of toric varieties): a new proof of the following theorem.
@@
Criterion 3.4.2 cannot be applied directly, since $\mathcal{C}$} contains the Tate twist and complex conjugation acts trivially on its characters of $\mathbb{G}_{m}$. We consider instead the category $\mathcal{C}$} of ungraded (i.e. defined by algebraic correspondences of any degree) $\lambda$-adic homological motives on powers of $A$, with coefficients in $K$. This does not affect the issue of numerical equivalence versus $\ell$-adic homological equivalence.
@@
Note. This new proof of the Clozel-Deligne theorem, which uses very few features from the geometry of abelian varieties, extends to categories $\mathcal{C}$} of ‘CM motives’, not related a priori to abelian varieties: those for which the ‘special numerical motivic group’ $G$ (the kernel of the map from the numerical motivic group to $\mathbb{G}_{m}$ associated to the subcategory of Tate motives in $\mathcal{C}$}) has the properties of Criterion 3.4.2.
Acknowledgements. The author thanks B. Kahn, H. Esnault and the referees for many suggestions of improvement of the presentation, and P. Deligne for suggesting to include a discussion of multifiltrations (§2.3.5).